# Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Cameras

### v2

@@ Abstract
The need for accurate and non-intrusive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid motion estimation. This study investigates the tremendous potential of spike cameras (a type of ultra-high-speed, high-dynamic-range camera) in PIV. We propose a deep learning framework, Spike Imaging Velocimetry (SIV), designed specifically for highly turbulent and intricate flow fields. To aggregate motion features from the spike stream while minimizing information loss, we incorporate a Detail-Preserving Hierarchical Transform (DPHT) module. Additionally, we introduce a Graph Encoder (GE) to extract contextual features from highly complex fluid flows. Furthermore, we present a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), which provides labeled data for three challenging fluid dynamics scenarios. Our proposed method achieves superior performance compared to existing baseline methods on PSSD. The datasets and our implementation of SIV are open-sourced in the supplementary materials.
1 Introduction
@@
The need for accurate and non-intrusive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid dynamics research. This technique has been widely applied to critical topics such as studying the aerodynamic shapes of aircraftKompenhans et al. (2000), optimizing vehicle designs to reduce air resistanceNakagawa et al. (2016), improving energy conversion efficiency in combustion processesChen et al. (2022), understanding the impact of blood flow dynamics on diseasesRaschi et al. (2012), and predicting the spread of water pollutionMoayeriKashani et al. (2016).
In PIV experiments or simulations, tracer particles(e.g., silver-coated particles) are introduced into the fluid to scatter optical signals. By estimating their displacements, PIV can infer the velocities of the fluid in which they are suspended.
@@
Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetSCai et al. (2019b) and RAFT-PIVLagemann et al. (2021).
@@
The success of optical flow networksDosovitskiy et al. (2015); Sun et al. (2018); Teed and Deng (2020) inspires the development of learning-based PIV algorithmsCai et al. (2019b, a); Lagemann et al. (2021). However, the performance improvements of these PIV networks mainly rely on advancements in optical flow estimation networks, rather than targeted improvements that incorporate the specific characteristics of the flow field. Although optical flow networks generally exhibit good generalization ability, the characteristics of fluid scenes cannot be ignored for further performance improvement. Turbulence is an irregular flow phenomenon occurring at high Reynolds numbers. When measuring the turbulent velocity field with PIV, this scenario is significantly different from macroscopic worlds, with the following key challenges:
@@
(2) Signals are only present in regions with tracer particles, and since the particle volume is small, downsampling may lead to signal loss;
(3) Unlike solids, the fluid is highly unstructured overall, despite the presence of vortices of different scales (called quasi-order structures).
These challenges often coexist in turbulence. Challenge (1) places high demands on experimental hardware, typically requiring ultra-high-speed cameras. Challenge (2) requires us to carefully address potential resolution loss in the network. Challenge (3) introduces problems to the extraction of context features. Based on these characteristics, we propose improvements in both hardware and algorithm aspects.
In terms of hardware, we explore the application of spike cameras (a type of neuromorphic camera) for PIV. We argue that using spike cameras instead of CMOS sensors in fluid measurement tasks offers several advantages:
High temporal resolution (HTR). Spike cameras can achieve ultra-high temporal resolution (20000Hz), reducing the interval $\Delta t$ between the source and target time, and providing rich temporal information for accurate predictions. This allows the spike camera to cope well with Challenge (1);
High dynamic range (HDR). When particles accumulate in large numbers, local regions may become overly bright, exceeding the dynamic range of conventional cameras. Spike cameras asynchronously accumulate photons at each pixel, inherently possessing a much larger dynamic range compared to CMOS sensors. This makes them more advantageous in scenarios with uneven brightness (such as inappropriate lighting or particles moving toward concentrated areas);
@@
In terms of algorithm, we introduce a specialized network, Spike Imaging Velocimetry (SIV) to estimate fluid displacement fields. We propose two modules, the Detail-Preserving Hierarchical Transform (DPHT) designed in consideration of Challenge (2), and the Graph Encoder (GE) mainly designed for Challenge (3). We also introduce a dataset called Particle Scenes with Spike and Displacement (PSSD). PSSD comprises labeled data addressing three common motion estimation scenarios: steady turbulence (Problem 1), high-speed flow (Problem 2), and high dynamic range scenarios (Problem 3). In the experimental section, we conduct comparative analyses of our method against both image-based and spike-based approaches. Our proposed SIV achieves state-of-the-art performance across all three sub-datasets. The effectiveness of the proposed modules is further demonstrated through ablation studies. The main contributions of this work are as follows:
@@
1. Explore a feasible approach for applying spike cameras to fluid velocity measurement;
2. From the perspective of fluid dynamics, especially turbulent flows, and the characteristics of the PIV scenarios, design a suitable network structure SIV, including two new modules: DPHT and GE;
@@ 1. Explore a feasible approach for applying spike cameras to fluid velocity measurement;
3. Propose a PIV dataset for recording continuous fluid motion with spike cameras, which includes four flow types and three challenging experimental scenarios.
2 Related Work
2.1 Spike-based Optical Flow Estimation
@@
The spike cameraDong et al. (2021); Huang (2019) is a type of neuromorphic camera (NeurCam). Inspired by biological visual neural structures and mechanisms, NeurCams emulate the encoding principles of the primate retina. Spike cameras mimic the encoding principles of the central fovea in the retina, providing continuous visual information with ultra-high temporal resolutions, high dynamic range(HDR), and significantly reduced energy consumption and latency, making them well-suited for high-speed motion scenarios. Their pixels independently accumulate photons and emit a spike when the accumulated intensity exceeds a certain threshold. This process maps the incoming photon stream into the spike stream, which preserves spatiotemporal relationships and can be reconstructed into images. For a detailed explanation of the working mechanism, refer to Dong et al. (2021). The application of spike cameras in optical flow estimation provides significant insights for this study. Although optical flow estimation is generally used to estimate the motion of macroscopic objects, its principles and objectives are aligned with those of Particle Image Velocimetry (PIV). The techniques employed in PIV for image registration and displacement estimation are essentially a variant of optical flow estimation. Spike-based optical flow estimation is still a relatively new field. Hu et al. (2022) introduced a specialized neural network architecture for spike-based optical flow estimation. This structure can adaptively eliminate motion blur in spike streams based on prior motion information. A framework based on Teed and Deng (2020) is proposed by Zhao et al. (2022) to handle continuous spike streams and is then extended by Zhao et al. (2024) to suppress the adverse effects of randomness in spike streams. Chen et al. (2023) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by Xia et al. (2024).
2.2 Image-based PIV Algorithms
@@
Many traditional image-based PIV algorithmsLourenco and Krothapalli (1995); Nobach and Honkanen (2005) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation areas (IA). Displacement is then determined by finding the maximum correlation between the two interrogation windows, which is similar to the correlation method in optical flow estimation to some extent. However, traditional methods depend on finely tuned parameters and complex optimization techniques. With advancements in optical flow networks, various architecturesDosovitskiy et al. (2015); Sun et al. (2018); Teed and Deng (2020) are developed to predict dense displacement fields of macroscopic objects with high accuracy. Early PIV networks like PIV-DCNNLee et al. (2017) laid the groundwork for these developments. PIV-NetSCai et al. (2019b) is capable of generating full-resolution displacement outputs. Cai et al. (2019a) adapt LiteFlowNetHui et al. (2018) to PIV. RAFT-PIVLagemann et al. (2021), derived from RAFTTeed and Deng (2020), has shown exceptional performance on multiple datasets. Additionally, unsupervised learning frameworks for PIVZhang and Piggott (2020); Zhang et al. (2023a) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements. Other recent PIV networks include LIMAManickathan et al. (2023), CC-FCNGao et al. (2021), En-FlowNetCJi et al. (2024a), ARAFT-FlowNetHan and Wang (2023), OFVNetSJi et al. (2024b), and PIV-PWCNetZhang et al. (2023b). However, the performance improvements of these PIV networks mainly rely on advancements in optical flow estimation networks, rather than targeted improvements that incorporate the specific characteristics of the flow field, and most of them have not been open-sourced.
2.3 Event-based PIV Algorithms
@@
The event camera(such as DVSLichtsteiner et al. (2008), DAVISMoeys et al. (2017), ATISPosch et al. (2010), CeleXHuang et al. (2017)) is another type of NeurCam, which is inspired by the peripheral retinaRebecq et al. (2019). Unlike spike cameras, event cameras independently monitor the change in light intensity (increase or decrease) at each pixel and generate signals based on the local pixel information of dynamic changes. When the brightness change of a pixel exceeds a preset threshold, the camera records this change event (brightness increase as a positive event, brightness decrease as a negative event) and outputs its timestamp, position, and change value as data. However, due to the lack of absolute intensity values, event cameras struggle to reconstruct complete images, particularly in static regions and with fine textures. The application of event cameras to PIV has already been explored by Willert and Klinner (2022) and Willert (2023). Similar to the spike camera, the event camera offers higher temporal resolution, high dynamic range, low latency, and reduced experimental complexity. However, event cameras are only sensitive to objects with obvious motion and have hardly any output signal in stationary regions. This can be a weakness compared to spike cameras, considering that not all areas of the flow field may be moving at high speed.
3 Preliminary
3.1 The Working Principle of the Spike Camera
@@
Where $A(\mathbf{x},t)$ represents the voltage of the accumulator at pixel $\mathbf{x}=(x,y)$ at time $\tau$, and $\theta$ is the threshold of the comparator. The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represent the time length of the spike stream. Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and brightness representation.
3.2 Data Generation
Our proposed dataset, Particle Scenes with Spike and Displacement (PSSD), includes three sub-datasets: Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3). Each sub-dataset comprises 10,000 sample sequences of spike streams and the corresponding ground truth turbulent displacement fields when $\Delta t=21$ (about 1.05 ms) and $\Delta t=11$ (about 0.55 ms). Additionally, we provide corresponding pairs of images for comparative analysis for image-based methods. We use the reference flow field to simulate the motion of the scene and construct a virtual spike camera to trigger the spike stream. The reference flow field data are taken from the high-precision numerical simulation database used for turbulence research, Johns Hopkins Turbulence Database (JHTDB)Li et al. (2008). JHTDB employs Direct Numerical Simulation (DNS) to simulate turbulent flow fields. DNS directly solves the Navier-Stokes equations without any model assumptions or simplifications, enabling precise capture of all scales in turbulent flow. This method resolves the fluid motion in both space and time with high-resolution grids, accurately capturing the complex and multi-scale nature of turbulence. Additional details about the dataset can be found in our supplementary materials.
4 Approaches
4.1 Overall Architecture
@@
Here, $\mathcal{I}$ is the mapping of the spike stream $\mathbf{S}$ in the high-dimensional feature space, representing the scene information, i.e., $\mathcal{I}=\mathrm{Rep}(\mathbf{S})$.
Network architecture. The primary structure of the SIV network is shown in Figure 2. To represent the scenes at moments $t$ and $t+\Delta t$, we extract spike stream segments (source and target) near these two moments, denoted as $\mathbf{S}_{s}$ and $\mathbf{S}_{t}$, and input them into the network. Firstly, through the proposed DPHT module, the sub-stream $\mathbf{S}_{s}$ and $\mathbf{S}_{t}$ are embedded into the representations $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$ by our proposed DPHT module. Motion features $\mathbf{F}_{s,m}$ and $\mathbf{F}_{t,m}$ are extracted from $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$. Through the proposed Graph Encoder, the context feature $\mathbf{F}_{c}$ is extracted from $\mathbf{R}_{t}$. Subsequently, the enhanced features $\mathbf{F}_{s,m}$, $\mathbf{F}_{t,m}$ and $\mathbf{F}_{c}$ undergo an iterative optimization process in a pyramid structure, ranging from coarse (1/16 downsampled branch) to fine (1/4 downsampled branch). In each branch, similar to DIPZheng et al. (2022), the Inverse Patchmatch Block is used to construct the correlation volume, and is aggregated through a GRU. Finally, the Local Search Block estimates the residual displacement field $\Delta\mathbf{D}$ by searching for local cost based on the current displacement field estimate, and cost aggregation is performed through another GRU.
Figure 2: The overall network structure of SIV includes motion feature extraction supported by DPHT and the motion encoder, context feature extraction supported by DPHT and the Graph Encoder, along with a multi-scale iterative optimization process.
Figure 3: (a) A demonstration of the Detail-Preserving Hierarchical Transform (DPHT) module. (b) A brief illustration of Graph Encoder (GE) module.
4.2 Detail-Preserving Hierarchical Transform
Zhao et al. (2024) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network. Low-level features mainly focus on local details, while high-level features aggregate these low-level features to express more complex meanings. However, our ablation experiments (see Table 2) show that directly using multi-level representations does not achieve optimal performance improvement for PIV. This is because, as mentioned in the Introduction, the reduction in resolution or scale of feature maps leads to a significant loss of detail information. Furthermore, these details may not be critical for macroscopic objects, but considering that the particles in PIV are small and the fluid motion is highly intricate and complex, even the loss of detail information could result in an increase in local prediction errors. Moreover, the sparse and dark regions without particles essentially carry little to no signal, much of the information in a small region may stem from the bright particles. Based on these considerations, we designed a 3D Detail-Preserving Downsampling unit in which a bunch of 3D convolutions and 3D max poolings are incorporated (as shown in Figure 3(a)) to reduce the progressive loss of important information across the multi-layer structure while enabling the network to aggregate features at multiple scales. We refer to the entire block in Figure 3(a) as the Detail-Preserving Hierarchical Transform (DPHT).
@@ 4.2 Detail-Preserving Hierarchical Transform
$\mathbf{R}=\mathscr{A}_{m}[\mathbf{x}^{(l=1)}_{\mathrm{out}}\oplus\mathbf{x}^{(l=2)}_{\mathrm{out}}\oplus\mathbf{x}^{(l=3)}_{\mathrm{out}}],$ (4)
$\oplus$ represents concatenation, and $\mathscr{A}_{m}$ denotes convolution. As shown in Figure 3, we use DPHT as a pre-representation method before motion feature extraction and context information extraction, obtaining $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$ for source sub-stream and target sub-tream, respectively.
4.3 Graph Encoder
Turbulent motion is highly complex, and extracting context information from turbulence is challenging. Therefore, Yu et al. (2021) argued that context extractors are unnecessary for PIV and removed this structure from their network. However, we hold a different view: despite the spatially chaotic nature of turbulence, there still exist multi-scale vortex structures (referred to as quasi-order structures), making the extraction of context information meaningful. To achieve this, we introduce Graph Convolutional Unit (GCU) in the context extractor, referring to the entire context extractor as the Graph Encoder. Graph convolutions weight the relationships between nodes using an adjacency matrix, performing local aggregation. Each node not only focuses on its own features but also passes information from neighboring nodes (as defined by the adjacency matrix) to itself. We believe that this approach is particularly suited for capturing spatial or topological structures in the PIV data, especially when the quasi-order structures of the flow is highly complex topologically. This situation arises particularly in cases with extremely high flow speeds (meaning high Reynolds numbers). When the Reynolds number is sufficiently high, turbulence experiences a cascade process from large-scale vortices to small-scale vortices (Kolmogorov scale) Kolmogorov (1991). This cascade process is the transfer of energy from large vortices to small vortices in turbulence, with each vortex disturbing the surrounding flow field, further complicating the flow.
@@ 4.3 Graph Encoder
Graph projection. The first step is to map the preliminary context information $\mathbf{R}_{c}=\mathbf{R}_{s}$ extracted from DPHT from the general coordinate space to the graph space. The projection operation decouples the position information from the original grid features, making the resulting low-dimensional node feature representations more compact and expressive. The mapped context node representations, denoted as $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}$, contain appearance features related to the shape and regional information of the scene context. Given $\mathbf{R}_{c}$, we can use the projection function $\mathscr{P}_{R\to V}(\cdot)$ to assign features with similar representations to the same node, resulting $\mathbf{V}_{c}\in\mathbb{R}^{C\times K}$, where $C$ indicates channel number and $K$ is a hyper-parameter representing the number of nodes. Since $\mathscr{P}_{R\to V}(\cdot)$ we define is a linear combination of feature vectors within the grid space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{c})$, the resulting nodes can theoretically aggregate both long-range and short-range information across the entire feature map, which could be beneficial for turbulent scenes with multi-scale quasi-order structures. This is formulated by,
$v_{c,i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{c,j}).$ (5)

### v3

# Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Streams

@@ Abstract
Particle Image Velocimetry (PIV) is a widely adopted non-invasive imaging technique that tracks the motion of tracer particles across image sequences to capture the velocity distribution of fluid flows. It is commonly employed to analyze complex flow structures and validate numerical simulations. This study explores the untapped potential of spike cameras—ultra-high-speed, high-dynamic-range vision sensors—in high-speed fluid velocimetry. We propose a deep learning framework, Spike Imaging Velocimetry (SIV), tailored for high-resolution fluid motion estimation. To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR). Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), comprising labeled samples from three representative scenarios in fluid dynamics. Our proposed method outperforms existing baselines on the PSSD dataset, demonstrating its effectiveness. The implementation details of our project, including code and configuration settings, are provided in the supplementary material.
Abstract
Particle Image Velocimetry (PIV) is a widely adopted non-invasive imaging technique that tracks the motion of tracer particles across image sequences to capture the velocity distribution of fluid flows. It is commonly employed to analyze complex flow structures and validate numerical simulations. This study explores the untapped potential of spike cameras—ultra-high-speed, high-dynamic-range vision sensors—in high-speed fluid velocimetry. We propose a deep learning framework, Spike Imaging Velocimetry (SIV), tailored for high-resolution fluid motion estimation. To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR). Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), comprising labeled samples from three representative scenarios in fluid dynamics. Our proposed method outperforms existing baselines on the PSSD dataset, demonstrating its effectiveness. The implementation details of our project, including code and configuration settings, are provided in the supplementary material.
Introduction
@@
The need for accurate and non-invasive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid dynamics research. This technique has been widely applied to critical topics such as studying the aerodynamic shapes of aircraft (Kompenhans et al. 2000), optimizing vehicle designs to reduce air resistance (Nakagawa et al. 2016), improving energy conversion efficiency in combustion processes (Chen, Li, and Zhu 2022), understanding the impact of blood flow dynamics on diseases (Raschi et al. 2012), and predicting the spread of water pollution (MoayeriKashani et al. 2016).
@@
Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetS (Cai et al. 2019b) and RAFT-PIV (Lagemann et al. 2021).
@@
In PIV experiments or simulations, tracer particles (e.g., silver-coated particles) are introduced into the fluid to scatter optical signals. By estimating their displacements, PIV can infer the velocities of the fluid in which they are suspended. The success of optical flow networks (Dosovitskiy et al. 2015; Sun et al. 2018; Teed and Deng 2020) has inspired the development of learning-based PIV algorithms (Cai et al. 2019b, a; Lagemann et al. 2021). Although optical flow networks generally exhibit good generalization ability, the characteristics of fluid scenes cannot be ignored for further performance improvement. Turbulence is an irregular flow phenomenon occurring at high Reynolds numbers. When measuring the turbulent velocity field with PIV, this scenario is significantly different from macroscopic worlds, with the following key challenges:
@@
(2) The scene contains extensive empty regions without tracer particles, yet the network is required to estimate the full flow field.
(3) Unlike solids, the fluid is highly sparse and unstructured overall, despite the presence of vortices of different scales (called quasi-order structures).
These challenges often coexist in turbulence. Challenge (1) places high demands on the experimental hardware, typically requiring ultra-high-speed cameras. Challenge (2) requires addressing potential resolution loss in the network carefully. Challenge (3) introduces problems to the extraction of context features. To address these challenges, we propose improvements in both hardware and algorithm aspects.
In terms of hardware, we explore the application of spike cameras (a type of neuromorphic camera) for PIV. We argue that using spike cameras in fluid measurement tasks offers several advantages:
High temporal resolution (HTR). Spike cameras can achieve ultra-high temporal resolution (20000Hz), reducing the interval $\Delta t$ between the source and target time, and providing rich temporal information for accurate predictions. This enables spike cameras to handle Challenge (1) effectively;
High dynamic range (HDR). When particles accumulate in large numbers, local regions may become overly bright, exceeding the dynamic range of conventional cameras. Spike cameras asynchronously accumulate photons at each pixel, inherently possessing a much larger dynamic range. This makes them more advantageous in scenarios with uneven brightness (such as inappropriate lighting or particles moving toward concentrated regions);
@@
We propose a specialized network, Spike Imaging Velocimetry (SIV), for estimating fluid velocity fields. To enhance network performance, we develop three novel modules tailored to the characteristics of fluid dynamics and spike cameras. These include the Detail-Preserving Hierarchical Transform (DPHT), which enhances feature extraction from spike streams while mitigating downsampling artifacts; the Graph Encoder (GE), which maps features into a graph structure for adaptive context aggregation; and the Multi-scale Velocity Refinement (MSVR), which addresses over-smoothing in predicted flow fields.
We also introduce a new dataset, Particle Scenes with Spike and Displacement (PSSD), designed for learning fluid motion from spike data. It covers three representative motion estimation scenarios: steady turbulence, high-speed flow, and high dynamic range conditions.
Extensive experiments demonstrate the superiority of SIV over both image-based and spike-based baselines. SIV achieves state-of-the-art performance across all evaluation settings. The effectiveness of each proposed module is further validated through ablation studies.
The main contributions of this work are summarized as follows:
1.
@@
We explore a feasible solution for applying spike cameras to fluid velocity measurement, opening new directions for PIV.
2. We design a dedicated network, SIV, tailored to fluid dynamics and turbulent flow properties, and propose three novel modules: the DPHT, GE, and MSVR.
@@
3. We propose a new dataset, PSSD, which captures continuous fluid motion across four flow types and three challenging experimental scenarios.
4. We conduct extensive quantitative experiments and qualitative visual comparisons, demonstrating that SIV achieves state-of-the-art performance across multiple benchmarks. We further perform ablation studies to systematically validate the effectiveness of each proposed module.
Related Work
Spike-based Optical Flow Estimation
@@
The spike camera (Dong, Huang, and Tian 2021; Huang 2019) is a type of neuromorphic camera (NeurCam). Inspired by biological visual neural structures and mechanisms, NeurCams emulate the encoding principles of the primate retina. Spike cameras mimic the encoding principles of the central fovea in the retina, providing continuous visual information with ultra-high temporal resolutions, high dynamic range(HDR), and significantly reduced energy consumption and latency, making them well-suited for high-speed motion scenarios. Their pixels independently accumulate photons and emit a spike when the accumulated intensity exceeds a certain threshold. This process maps the incoming photon stream into the spike stream, which preserves spatiotemporal relationships and can be reconstructed into images. For a detailed explanation of the working mechanism, refer to (Dong, Huang, and Tian 2021). The application of spike cameras in optical flow estimation provides significant insights for this study. Although optical flow estimation is generally used to estimate the motion of macroscopic objects, its principles and objectives are aligned with those of PIV. The techniques employed in PIV for image registration and displacement estimation are essentially a variant of optical flow estimation. Spike-based optical flow estimation is still a relatively new field. (Hu et al. 2022) introduced a specialized neural network architecture for spike-based optical flow estimation. This structure can adaptively eliminate motion blur in spike streams based on prior motion information. A framework based on RAFT is proposed by (Zhao et al. 2022) to handle continuous spike streams. (Zhao et al. 2024) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network. (Chen, Yu, and Huang 2023) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by (Xia et al. 2024).
Image-based PIV Algorithms
@@
Many traditional image-based PIV algorithms (Lourenco and Krothapalli 1995; Nobach and Honkanen 2005) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation windows. Displacement is then determined by finding the maximum correlation between the two interrogation windows, which is similar to the correlation method in optical flow estimation to some extent. However, traditional methods depend on finely tuned parameters and complex optimization techniques. With advancements in optical flow networks, various architectures (Dosovitskiy et al. 2015; Sun et al. 2018; Teed and Deng 2020) have been developed to predict dense velocity fields of macroscopic objects with high accuracy. Early PIV networks like PIV-DCNN (Lee, Yang, and Yin 2017) laid the groundwork for these developments. PIV-NetS (Cai et al. 2019b) is capable of generating full-resolution displacement outputs. (Cai et al. 2019a) adapt LiteFlowNet (Hui, Tang, and Loy 2018) to PIV. RAFT-PIV (Lagemann et al. 2021), derived from RAFT (Teed and Deng 2020), has shown exceptional performance on multiple datasets. (Yu et al. 2021) removes the computationally inefficient components of RAFT, such as the context encoder, to improve inference speed. Additionally, unsupervised learning frameworks for PIV(Zhang and Piggott 2020; Zhang et al. 2023a) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements. Other recent PIV networks include LIMA (Manickathan, Mucignat, and Lunati 2023), CC-FCN (Gao et al. 2021), En-FlowNetC (Ji, An, and Hui 2024), ARAFT-FlowNet (Han and Wang 2023), OFVNetS (Ji, Hui, and An 2024), and PIV-PWCNet (Zhang et al. 2023b). However, the performance gains of most PIV networks are largely attributed to advances in optical flow estimation architectures, rather than targeted enhancements tailored to the specific characteristics of fluid dynamics, and most implementations remain unavailable as open-source.
Event-based PIV Algorithms
@@
The event camera (Lichtsteiner, Posch, and Delbruck 2008; Moeys et al. 2017; Posch, Matolin, and Wohlgenannt 2010; Huang, Guo, and Chen 2017) is another type of NeurCam, which is inspired by the peripheral retina (Rebecq et al. 2019). Unlike spike cameras, event cameras independently monitor the change in light intensity (increase or decrease) at each pixel and generate signals based on the local pixel information of dynamic changes. When the brightness change of a pixel exceeds a preset threshold, the camera records this change event (brightness increase as a positive event, brightness decrease as a negative event) and outputs its timestamp, position, and change value as data. However, due to the lack of absolute intensity values, event cameras struggle to reconstruct complete images, particularly in static regions and with fine textures. The application of event cameras to PIV has already been explored by (Willert and Klinner 2022) and (Willert 2023). Similar to the spike camera, the event camera offers higher temporal resolution, high dynamic range, low latency, and reduced experimental complexity.
Preliminary
The Working Principle of the Spike Camera
@@
Where $A(\mathbf{x},t)$ represents the voltage of the accumulator at pixel $\mathbf{x}=(x,y)$ at time $\tau$, and $\theta$ is the threshold of the comparator. The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represents the temporal dimension. Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and brightness representation.
Data Generation
Our proposed dataset, Particle Scenes with Spike and Displacement (PSSD), includes three sub-datasets: Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3). Each sub-dataset comprises 10,000 sample sequences of spike streams and the corresponding ground-truth turbulent velocity fields when $\Delta t=21$ (approximately 1.05 ms) and $\Delta t=11$ (approximately 0.55 ms). Additionally, we provide corresponding pairs of images for comparative analysis for image-based methods. We use the reference flow field to simulate the motion of the scene and construct a virtual spike camera to trigger the spike stream. The reference flow field data are taken from the high-precision numerical simulation database used for turbulence research, Johns Hopkins Turbulence Database (JHTDB) (Li et al. 2008). JHTDB employs Direct Numerical Simulation (DNS) to simulate turbulent flow fields. DNS directly solves the Navier-Stokes equations without any model assumptions or simplifications, enabling precise capture of all scales in turbulent flow. This method resolves the fluid motion in both space and time with high-resolution grids, accurately capturing the complex and multi-scale nature of turbulence. Additional details about the dataset can be found in our supplementary materials.
Approaches
Overall Architecture
@@
Here, $\mathcal{I}$ is the mapping of the spike stream $\mathbf{S}$ in the high-dimensional feature space, representing the scene information.
Figure 2: Spike Imaging Velocimetry (SIV) network.
Figure 3: (a) Detail-Preserving Hierarchical Transform (DPHT) Module. (b) Graph Encoder (GE).
Detail-Preserving Hierarchical Transform
In order to avoid the loss of particle signal information during feature extraction, we propose a multi-scale pyramidal architecture, named Detail-Preserving Hierarchical Transform (DPHT), which incorporates a 3D Detail-Preserving Downsampling module (as shown in Figure 3(a)).
@@
$\mathbf{R}=\mathscr{A}_{m}[\mathbf{Concat}(\mathbf{x}^{(l=1)}_{\mathrm{out}},\mathbf{x}^{(l=2)}_{\mathrm{out}},\mathbf{x}^{(l=3)}_{\mathrm{out}})$], (4)
$\mathscr{A}_{m}$ denotes convolution. After processing with DPHT, the signals of the source and target sub-streams are denoted as $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$, respectively.
Graph Encoder
When the Reynolds number is high, fluid flow becomes turbulent. As energy cascades and dissipates, large vortices become unstable and gradually break into smaller ones until they vanish. This indicates that fluid motion is not purely chaotic, but rather exhibits an underlying topological structure. To capture the contextual information of the fluid, we propose a Graph Encoder (GE). It is a feature extraction network based on Graph Neural Networks (GNNs) and Convolutional Neural Networks (CNNs). By representing features as graph nodes and constructing a dynamic adjacency matrix, GE flexibly aggregates and updates information from neighboring nodes, enabling a larger receptive field than conventional CNNs. Graph Attention Networks (GATs) (Veličković et al. 2017), leveraging attention mechanisms, learn the relative contribution of each neighbor to a node’s new representation, thereby mitigating interference from particle-sparse regions. The serial combination of GNN and CNN enables adaptive global perception and local refinement, effectively capturing fluid-specific contextual features.
@@
Graph projection. The first step is to map $\mathbf{R}_{s}$ extracted from DPHT from the general coordinate space to the node space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{s})$. The mapped context node representations, denoted as $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}\in\mathbb{R}^{C\times K}$, contain appearance features related to the shape and regional information of the scene. $C$ indicates channel number and $K$ is a hyper-parameter representing the number of nodes. In this paper, we set $K=128$. $\mathscr{P}_{R\to V}(\cdot)$ is a linear combination of feature vectors within the grid space, formulated by,
$v_{c,i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{s,j}).$ (5)